LControl

LANC/Control-L Interface for Amiga Computers Version 1.41 March 1995

by Juergen Frank und Michael Watzl

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993-95 Juergen Frank und Michael Watzl

1 What is LControl

LControl is a program to control video cameras via a special control hardware. The cameras must have a control-L or LANC interface (Sony kompatible).

The hardware is very easy to construct. Schematics, construction kit or the complete hardware will be sendt to you after registering, See Appendix B [Registering], page 8.

The control hardware is connected to the parralel port. Two cameras can be connected to the hardware. You can switch between the two cameras via the software.

LControl is a commodity and it has an arexx port. Therefore LControl can work in the background and your camera can be "programmed" via arexx.

See also Appendix C [IR-Master-Support], page 9.

2 Installation

Requirements:

• Kickstart 2.04 or better

• Workbench 2.04 or better

1. Copy ‘LControl’ into any drawer you want.

2. Copy ‘LANC.library’ to ‘LIBS:’

3. If you use WB2.1 or better: You do not need to install an english catalog - it is built in. For other languages (so far, only german is supported):

Copy the file ‘lcontrol.catalog’ from catalogs/<your language> to LOCALE:catalogs/<your language>.

When you copy LControl to the ‘WBStartup’-drawer, you should also add the DONOTWAIT tooltype.

3 How To Use

3.1 Menus

3.1.1 Project

About... Some information about the authors and the program Hide Closes the LControl window. However LControl stays active. Fix position

Snapshots the current position. This piece of information is saved in the tooltypes, See Section 3.4 [Tooltypes], page 6.

Quit Quits LControl

3.1.2 Special

Reset counter

Resets the counter of the camera.

Zoom in Sends "Zoom In" to the camera Zoom out Sends "Zoom Out" to the camera Record Sends "Record" to the camera. Eject Sends "Eject" to the camera. Gotomode

In the submenu you can select one of these commands.

0 Play 1 Stop 2 Pause 3 Record The selected command is sent when a successful "goto" was made.

Goto You may enter a counter. Then the tape is winded to this position an the command selected with gotomode is sent.

3.1.3 Special-II

   Note:

All the commands in this menu refer directly to standard camera commands. Please look up their meaning in your camera’s user manual.

3.2 Gadgets

Swap Swaps to the other camera Play Sends a "Play"-command to the camera Stop Sends a "Stop"-command to the camera Pause Sends a "Pause"-command to the camera FREW Sends a "fast rewind" command to the camera FFOR Sends a "fast forward" to the camera PREW Sends a "picture rewind" to the camera PFOR Sends a "picture forward" to the camera −− Sends a "slow motion rewind" to the camera − Sends a "single pic back" to the camera + Sends a "singse pic forward" to the camera ++ Sends a "slow motion forward" to the camera

3.3 ARexx Port

The name of the arexx port is ‘LControl_rexx’

So far, the following commands are supported:

‘LC_QUIT’ Quits LControl ‘LC_SHOW’ Opens the LControl window ‘LC_HIDE’ Closese the LControl window

‘LC_ENABLE’

Opens the LControl window and opens the arexx port (if closed)

‘LC_DISABLE’

Closes the LControl window and the arexx port

‘LC_PLAY’
‘LC_STOP’
‘LC_PAUSE’
‘LC_FREW’
‘LC_FFOR’
‘LC_PREW’
‘LC_PFOR’
‘LC_PLUS’
‘LC_MINUS’
‘LC_SLOW’
‘LC_RSLOW’
‘LC_GETCOUNTER’

the current counter will be returned in RESULT

‘LC_GETSTATUS’

returns the current status in RESULT:

NO CAMERA

If no camera/hardware is connected

CAMERA0: <action>

Slot 1: active camera is connected.

CAMERA1: <action>

Slot 2: active camera is connected.

‘LC_RESET’

resets the counter of the camera

‘LC_SWAP’ swaps to the other camera

‘LC_ZOOMIN’

Zooms in one step.

‘LC_ZOOMOUT’

Zooms out one step.

‘LC_RECORD’

Sends Record to the camera.

‘LC_EJECT’

Sends Eject to the camera.

‘LC_GOTOMODE’

Expects a number (0,1,2 or 3) as parameter. See Section 3.1.3 [Menus], page 3.

‘LC_GOTO’ Expects a counter as parameter. See Section 3.1.3 [Menus], page 3.

‘LC_INSAUDIO’
‘LC_INSVIDEO’
‘LC_ASSEMBLE’
‘LC_INSA_LON’
‘LC_INSA_LOFF’
‘LC_INSA_RON’
‘LC_INSA_ROFF’
‘LC_INSV_ON’
‘LC_INSV_OFF’
‘LC_ASSEMBLEON’
‘LC_ASSEMBLEOFF’
‘LC_RAWCOMMAND’

This command directly sends data to the camera.

WARNING:Use this command only when you really know what you are doing!

Expects hexadezimal bytes as parameters. Example:

      LC_Rawcommand ff 10 aa

This will send 255, 16 and 170 to your camera.

Basic VTR-commands
hex                 command
00                  CH-1/1
40                  CH-2/2
20                  CH-3/3
60                  CH-4/4
10                  CH-5/5
50                  CH-6/6
30                  CH-7/7
70                  CH-8/8
08                  CH-9/9
48                  CH-10/10
28                  CH-11
68                  CH-12/CH/ENTER/#
18                  CH-13/1-
58                  CH-14/2-
38                  CH-15
78                  CH-16
04                  CH-HIGH
44                  CH-LOW
14                  X3 OR 2X
54                  POWER ON/OFF

22                  SLOW 2
62                  SLOW 3
32                  HIGH DOUBLE SPEED
2A                  ANT-SW
5A                  ANT-VTR
3A                  POWER ON
7A                  POWER OFF
06                  REVERSE
46                  FORWARD
09                  INDEX WRITE
49                  INDEX ERASE
65                  INDEX

3.4 Tooltypes

The following tooltypes are recognized:

CX POPUP= "YES" or "NO"

YES Pop window up on startup

NO Start hidden

CX POPKEY= Hotkey for LControl E.g.: lshift f10 ↦ left shift-key + F10

 Default: lcommand esc.

LEFT=xxx Leftedge of the window

TOP=yyy Topedge of the window

PUBSCREEN=name If specified, LControl tries to open its gui on the pbulicscreen named ‘name’.

GOTOMODE= Expects a decimal number (0,1,2 or 3) as parameter. See Section 3.1.3 [Menus], page 3, gotomode. You can specify the default command which should be sent when positioning with the goto-function.

LIBNAME= This should always be set to "LANC". This is for further expansion (Rapid, Panasonic).

Appendix A Support & Bugreports

A.1 Support

You may simply send a disk and enough stamps to one of the following addresses to get the newest software version of LControl:

Michael Watzl               or      Jürgen Frank
Konradstr. 11                       Wittelsbacherweg 7
86609 Donauwörth                    86609 Donauwörth
Tel.: +49 906 5834                  Tel.: +49 906 1057
Germany                             Germany

A.2 Bugreports

Bugreports or suggestions are welcome and should be sendt to one of the addresses above mentioned.

Please include a brief description of your computer and its configuration — this will help us to reproduce the bugs.

Appendix B Registering

REGISTRATION FORM (1.41)
NAME:       ___________________________________________
STREET:     ___________________________________________
ZIP-CODE:   ______CITY:     ________________________
COUNTRY:    ____________________TEL: _________________
I want to become registered user of LControl, therefore I
o enclose a cheque
  (europe: euro cheque
   others: drawn on a german bank)
o transfer payment to
Raiffeisen-Volksbank Donauwoerth eG
86609 Donauwoerth
Acc: 4185455   Bank code: 722 901 00
Comment"LCONTROL REGISTRIERUNG"
Please send me:
o Latest version + scematics for the hardware
  for $12 / DM 20
o Latest version + contruction kit
  for $18 / DM 30
o Latest version + complete Hardware
  for $25 / DM 40
Optional Hardware:
[for the construction kit]
o Two camera set [allows to connect two cameras and
  switch between them]
  plus $3 / DM 5
[for complete hardware]
o Two camera set [allows to connect two cameras and
  switch between them]
  plus $6 / DM 10
_______________________________________________
locality, date                            sign

Appendix C IR-Master-Support

Taken from "IR-Master.guide":

"The IR-Master project consists of two software components and a control hardware, which is connected to the joystick port of your Amiga.

At first you start editing your virtual remote control with the IR-Editor… This one is saved to disk then. Afterwards you activate the IR-Runner…

Your edited remote control opens as a window on the workbench. With this window you can interactively control your device[s], you can use the IR-Runner’s ARexx-port for programming your devices… Finally you can send infra red codes from within other applications or edit scripts which will automatically control e.g. your video recorder.

Since version 2.5 of the IR-Master two new features were introduced: You can directly send commands to LControl and Rapid. This allows you to control video cameras supported with an LANC/Control-Linterface and video recorders supported with an

GSE-Rapid-interface. . ."

Appendix D History

V1.00
    first Release
V1.01
    - arexx-command: GETSTATUS added
    - fixed some minor bugs
V1.02
    - Support for 2 cameras added [Switch]
    - minor bugfixes
V1.05
    - Zoom in/Zoom out added
    - Record added
V1.1
    - Goto added
    - Better SWAP implementation
V1.2
    [never released]

V1.3
- Now LControl uses a library (LANC.library)
- goto added (automatic positioning)
- many other minor bugfixes an features
V1.4
- localized

Index 11

Index

-

-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

– . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

+

+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 ++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

A

About... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 ARexx Port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Authors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

B

Bugreports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

C

Control-L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 CX POPKEY= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 CX POPUP= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

E

Eject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

F

FFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Fix position. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 FREW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

G

Goto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Gotomode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 GOTOMODE=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

H

Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Hide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 How to use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

I

Infra Red . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 IR-Master-Support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

K

Kickstart . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

L

LANC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 LC ASSEMBLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC ASSEMBLEOFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC ASSEMBLEON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC DISABLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC EJECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC ENABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC FFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC FREW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC GETCOUNTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC GETSTATUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC GOTO. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC GOTOMODE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC HIDE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 LC INSA LOFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSA LON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSA ROFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSA RON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSAUDIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSV OFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSV ON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSVIDEO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC MINUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PAUSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PLAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PLUS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PREW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC QUIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 LC RAWCOMMAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Index 12

LC RECORD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC RESET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC RSLOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC SHOW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 LC SLOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC STOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC SWAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC ZOOMIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC ZOOMOUT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LEFT=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 LIBNAME= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

M

Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

P

Parallel port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Pause. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 PFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Play . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 PREW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Prices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 PUBSCREEN=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Q

Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

R

Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Registering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Reset counter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

S

Sony. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Special . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Special-II . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Swap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

T

Tooltypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 TOP= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

U

Update . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

V

Videocamera. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

W

What is LControl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Z

Zoom in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Zoom out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Table of Contents

1 What is LControl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 3 How To Use. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3.1 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.1.1 Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.1.2 Special. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.1.3 Special-II . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.2 Gadgets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.3 ARexx Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.4 Tooltypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Appendix A Support & Bugreports. . . . . . . . . . . . . . . . . . . 7

A.1 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 A.2 Bugreports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

Appendix B Registering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Appendix C IR-Master-Support . . . . . . . . . . . . . . . . . . . . . . 9 Appendix D History. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11